# Robust Hallucination Detection in LLMs via Adaptive Token Selection

### v1

@@ 4.2 Augmenting Internal State Representations with Predictive Uncertainty
A lot of studies have demonstrated that uncertainty measurements are effective for hallucination detection. Our approach seeks to determine whether incorporating uncertainty metrics into the original token representation space can augment these representations’ ability to distinguish truthfulness.
The collection of predictive uncertainty measurements can be categorised into three levels: i) token-level uncertainty, i.e., predictive logits:
$p_{i}=P(y|y_{<i},\mathbf{x}),$ (7)
where $\mathbf{x}$ are prompt tokens and $y$ are generated tokens; ii) sentence-level perplexity, which is monotonically related to the mean of the negative log-likelihood of the sequence:
$P_{s}=-\frac{1}{T}\sum_{t=1}^{T}\log P(y|y_{<i},\mathbf{x}),$ (8)
and iii) semantic consistency across multiple samples, where the uncertainty value can be quantified by the number of semantic-equivalence generations over the whole generations based on the entailment results Farquhar et al. (2024). Specifically,
$P_{c_{m}}=\sum_{1}^{M}\frac{I_{c}=c_{m}}{M},$ (9)
where $M$ is the total number of generations, $c_{m}$ is the identified cluster and $I_{c}$ is the assigned cluster identity of a given sample. These uncertainty metrics can be directly leveraged for hallucination detection, and on average, the semantic consistency outperforms the other two metrics. However, it requires more computational cost while the other two metrics can obtained from just one generation as long as we can access the internal states of LLMs.
To augment the internal representations with the uncertainty information, we define the final input representation for each token as follows:
$\mathbf{h}^{\prime}=(\lambda_{1}+\lambda_{2}\cdot P_{uncertainty})\cdot\mathbf{h},$ (10)
where $\lambda_{1}$ and $\lambda_{2}$ are used to control the impact of uncertainty metrics. The improvements gained by various uncertainty measurements are evaluated and discussed in Section 5.4.
LLaMA-2-chat-7B LLaMA-2-chat-13B
Trivia QA SQuAD NQ BioASQ Trivia QA SQuAD NQ BioASQ
SE Farquhar et al. (2024) 0.879 0.799 0.801 0.823 0.777 0.767 0.769 0.743
p(True) Kadavath et al. (2022) 0.644 0.609 0.533 0.569 0.580 0.579 0.565 0.612
Perplexity Ren et al. (2023) 0.747 0.634 0.683 0.594 0.758 0.683 0.691 0.707
RMD Ren et al. (2023) 0.531 0.525 0.555 0.603 0.530 0.543 0.519 0.566
HaloScope Du et al. (2024) 0.625 0.574 0.615 0.682 0.564 0.543 0.575 0.624
LP-First Li et al. (2024) 0.796 0.760 0.715 0.823 0.690 0.699 0.670 0.720
LP-Last Kossen et al. (2024) 0.826 0.755 0.741 0.739 0.689 0.724 0.745 0.639
HaMI (Ours) 0.923 0.812 0.823 0.845 0.839 0.783 0.778 0.792
Table 1: AUCROC results of HaMI and seven SOTA competing hallucination detectors on four datasets with two LLaMA-based LLMs of different size. The results with Mistral are shown in the supplementary due to space limitation.
@@ 5 Experiments
We evaluate HaMI across various datasets and models. In this section, we first illustrate the experimental setup in Section 5.1 and present the main results in Section 5.2. We also conduct in-depth studies to analyse the robustness of our approach in Section 5.3 and the effectiveness of various components in Section 5.4.
5.1 Setup
@@
We investigate our methods on four question-answering (QA) datasets across a range of domains, including (1) Trivia QA, a relatively complicated confabulation QA datasets revealed by Joshi et al. (2017); (2) Stanford Question Answering Dataset (SQuAD for short) Rajpurkar (2016), which is based on Wikipedia and generated by humans through crowdsourcing; (3) Natural Questions (denoted as NQ) Kwiatkowski et al. (2019), containing search information from real users on Google search and (4) a biomedical QA corpus BioASQ Krithara et al. (2023). For each dataset, we randomly extract $3,000$ QA pairs for training and $800$ pairs for testing. Consistent with multi-sampling approaches, we prompt LLMs six times for each question for testing data. Following Farquhar et al. (2024), we have $400$ questions randomly sampled from the generated test set at the testing stage for evaluation.
[…]
@@
We compare our method with a series of baselines, covering uncertainty-based methods and internal representation-based methods. The uncertainty-based methods are as follows: Semantic Entropy (SE), semantic-equivalence measurement across multiple samples Farquhar et al. (2024), which employ GPT-3.5 for consistency evaluation; p(True) Kadavath et al. (2022), asking LLMs to express the probability or correctness for given answers themselves; Perplexity Ren et al. (2023), which is based on the Eq. 8. For internal representation-based methods, we choose the method based on relative Mahalanobis distance (RMD) Ren et al. (2023) and probing classifiers with different settings on supervising signals and token locations. We employ HaloScope Du et al. (2024), which proposes an automated membership estimation score and converts the score to binary labels for classification. For various location selection strategies, we choose the first and the last token as settled in many works, denoted as LP-First and LP-Last respectively Li et al. (2024); Orgad et al. (2024); Azaria and Mitchell (2023); Marks and Tegmark (2024); Kossen et al. (2024).
@@
Following Farquhar et al. (2024); Ren et al. (2023), we evaluate the model’s capability for hallucination detection by calculating the area under the receiver operating characteristic curve (AUROC), which is a widely used metric to evaluate the discriminative ability of binary classification. The ground-truth labels indicating the correctness of answers are given by GPT-4 Achiam et al. (2023), which is prompted to determine if the answer is correct or not based on the consistency between generated answers and gold answers and their own knowledge. The prompt template refers to Farquhar et al. (2024). Note that since GPT-4 makes mistakes mostly for the positive samples as checked with the gold answers, we ask GPT-4 to rejudge samples labelled as positive and discard samples if the result is inconsistent with the first result.
@@ 5.2 Main Results
In Table 1, we evaluate our proposed HaMI by comparing it with seven competing SOTA detection methods on four diverse QA datasets in LLMs of two different sizes. As depicted in the table, our approach achieves superior performance compared with other methods in all cases. In particular, HaMI outperforms others by a large margin on the Trivia QA dataset in both LLaMA-2-chat-7B and LLaMA-2-chat-13B models. Notably, both p(True) and Semantic Entropy resort to external LLMs for assistance, but their capability for hallucination detection differs significantly. Following the configurations described in their respective original studies, p(True) employs the LLaMA-7B model, while Semantic Entropy harnesses the more powerful GPT-3.5 model. This variation underscores that the performance of uncertainty-based methods, which rely on external support, is highly dependent upon the capabilities of the selected assistant LLM. Furthermore, among internal representation-based methods, supervised approaches generally outperform the unsupervised ones. For example, LP-First, LP-Last and our approach exhibit superior performance than HaloScope. LP-First can achieve impressive performance on some datasets like BioASQ but fail to work well on the other datasets. On the other hand, LP-Last works well on datasets like Trivia QA and less effectively on the other datasets. These results showcases the unstable performance of predetermined token-based methods. By contrast, our proposed HaMI performs consistently well on all four benchmarks, surpassing all competing methods. This superiority is due to not only the predictive uncertainty enhancement but also the effectiveness of adaptive token selection on generation responses of diverse forms (see Section 5.4 for more detailed analysis).
@@
Figure 3: AUROC results of cross-dataset generalisation on four datasets. Experiments are based on LLaMA-2-chat-7B.
@@ 5.3 Cross-dataset Generalisation Ability
The ability to generalise across datasets is essential to facilitating real-world applications of LLMs in diverse domains. We conduct experiments on the aforementioned four datasets to assess whether the proposed HaMI can effectively generalise among various datasets. For each dataset, we report the average AUROC scores of detectors trained on one of the other three datasets.
The cross-dataset generalisation performance is illustrated in Figure 3. It is clear that our method HaMI achieves consistently the best generalisation performance across all four datasets, outperforming the best competing method LP-First by 8% - 18%. Compared to the within-dataset performance in Table 1, the maximum performance decline in HaMI is no more than $4.0\%$, observed on the Trivia QA dataset, which is significantly lower than the competing methods. Notably, when trained on the other datasets, HaMI achieves an average AUROC score of $0.819$ on the SQuAD dataset, which marginally exceeds the score of $0.813$ when directly leveraging the SQuAD dataset as the training set. These findings affirm the effectiveness of HaMI as a reliable hallucination detector, demonstrating its adaptability in varied contexts.
@@ 5.4 Ablation Study
In this section, we comprehensively evaluate the effectiveness of different modules of our proposed HaMI.
Effectiveness of different modules.
In Table 2, we offer a detailed comparison of the improvements provided by various modules. The Baseline model is the one that uses the original internal representation of the First token as model inputs and does not use any of our modules. For adaptive token selection (ATS), we further explore the importance of the two loss functions, including both the MIL loss and the smoothness loss. Our results indicate that both of our ATS module and uncertainty-based semantic augmentation respectively contribute significant significant improvement over the baseline. In comparison to the augmentation module, the improvement from the ATS module is much larger. Within ATS, the smoothness loss can result in additional detection enhancement. Moreover, apart from the BioASQ dataset, substantial improvements are observed with the implementation of ATS following the integration of the semantic consistency score. This finding is especially crucial for practical applications, as the ATS module does not introduce additional computational burdens.
$P_{c_{m}}$ ATS TQ SQ NQ BQ
$\mathcal{L}_{MIL}$ $\mathcal{L}_{Smooth}$
0.799 0.714 0.656 0.780
0.858 0.756 0.724 0.828
0.858 0.785 0.795 0.803
0.906 0.787 0.776 0.839
0.923 0.812 0.823 0.845
Table 2: Experimental results on the effectiveness of the proposed uncertainty-augmented internal representations and the ATP module. Trivia QA, SQuAD, and BioASQ are denoted as TQ, SQ, and BQ for presentation respectively. Experiments are based on LLaMA-2-chat-7B.
Location Trivia QA SQuAD NQ BioASQ
First 0.858 0.756 0.724 0.828
Last 0.874 0.768 0.788 0.784
Mean 0.900 0.806 0.805 0.836
Ours 0.923 0.812 0.823 0.845
Table 3: AUROC results on the analysis of our adaptive token selection module. Experiments are based on LLaMA-2-chat-7B.
@@
Figure 4: Adaptive token selection results showing the top two tokens with the largest hallucination scores.
In-depth analysis of adaptive token selection.
Given that hallucinations may appear occasionally in long-form generations, identifying the critical token is important for capturing sufficient truthful information. Our proposed adaptive token selection (ATS) module solves the above concern. We conduct experiments to investigate the performance of the proposed ATS module, comparing it against commonly used benchmarks such as the First generated token, the Last generated token, and the Mean of all tokens. The results in Table 3 demonstrate that our selection strategy outperforms the alternatives across all datasets, yielding an average improvement of $8\%$ over the First token and $4\%$ over the Last token respectively. Compared with the single selection strategy, the Mean token can achieve more significant performance. We notice that the First token achieves better performance on the BioASQ dataset while showing a compromised behaviour on the other three datasets. The Last token is just the opposite. Considering that the average number of generated tokens for each question of the BioASQ dataset (i.e., $56$) is larger than the left datasets (Trivia QA - $22$, SQuAD - $30$, NQ - $30$), it is concluded that for longer generation, the last token may fail to retain important truthful information captured by earlier parts of the sequence. Conversely, in shorter generations, the last token can capture full semantics while the first token may miss content in the subsequent predictions. This suggests that i) the performance of the previous predefined token location-based approach could be sensitive to the length of the text generated and ii) our ATS module can well mitigate this issue.
@@
Figure 4 presents an illustrative example of the scoring results over tokens in a positive bag and a negative bag, where we can observe that there is significant distinguishability between the positive and negative tokens since the maximum scores of the instances in the positive bag is substantially greater than that in the negative bag. Tokens denoted by blue characters have scores that are very close to the largest ones and we find that these tokens can be adaptively identified by our method as concentrated answers in comparison to ground-truth answers. Additionally, it is noted that while selected tokens are associated with the exact answer, they may appear at any location in its vicinity, which can be captured by our smoothness loss.
Metrics Trivia QA SQuAD NQ BioASQ
Original $h_{i}^{l}$ 0.858 0.785 0.795 0.803
$\;\;+p_{i}$ 0.867 0.802 0.800 0.823
$\;\;+P_{s}$ 0.883 0.798 0.805 0.825
$\;\;+P_{c_{m}}$ 0.923 0.812 0.823 0.845
Table 4: AUROC results on the effectiveness of different uncertainty quantified metrics. $p_{i}$, $P_{s}$, and $P_{c_{m}}$ refer to logits, perplexity and semantic consistency respectively. Experiments are based on LLaMA-2-chat-7B.
Analysis of different uncertainty enhancement methods.
@@
As shown in Table 2, integrating uncertainty metrics into the original representation space can enhance the distinguishability between correct and hallucinated generations. There are different methods for this integration, as discussed in Section 4.2. Here we systematically examine using different uncertainty measuring methods for the internal representation enhancement, including token-level logit $p_{i}$, sequence-level perplexity $P_{s}$, and semantic consistency $P_{c_{m}}$ across multiple samples, as detailed in Eq. 7, 8 and 9 respectively. The final inputs are derived using Eq. 10. Experimental results are reported in Table 4, where we can observe that all three uncertainty metrics contribute to improved detection capabilities over the baseline representations. Specifically, the enhancements attributed to semantic consistency across multiple generations are particularly significant, exhibiting improvements up to $8.3\%$. Although the improvements observed with the other two metrics are less pronounced, both methods surpass the performance of the commonly used binary classifier using original representations. Moreover, we notice that incorporating sequence-level uncertainty enables our method to achieve performance comparable to the SOTA multi-sampling approach, namely Semantic Entropy (SE), without incurring the costs associated with multiple generations and external LLM employments. This observation highlights the potential of HaMI for deployments in various practical environments that involve external tools or not.
Figure 5: Impact of different layers on hallucination detection. The results are based on LLaMA-2-chat-7B. The suffix ‘-ori’ denotes experimental results with original representations.
How does HaMI perform with representations from different layers?
@@
We evaluate how layers impact the performance of HaML for hallucination detection with representations extracted from all 32 layers of the LLaMA-2-chat-7B model. Our investigation is based on the Trivia QA and NQ datasets, applying both original internal representations and semantic-equivalence-enhanced representations. As depicted in Figure 5, the AUROC values for using the original representations (indicated by the dashed line) exhibit a clear increase, peaking between layers 12 and 18, before declining to a relatively stable level between $0.75$ to $0.80$. This observation suggests that the truthfulness content evolves across the initial to middle layers. The performance of the uncertainty-enhanced representations, illustrated by solid lines, maintains a relatively consistent trend, with the highest AUROC scores concentrated in the middle layers. The comparison of these trends indicates that incorporating predictive uncertainty enhances the distinctiveness of the representations, particularly in the earlier layers where less semantic information is typically available.

### v2

@@ 4.2 Augmenting Internal State Representations with Predictive Uncertainty
Many studies have demonstrated that both uncertainty measurements and internal representations can capture truthfulness-related information, motivating us to explore the synergy of them to enhance hallucination detection. Moreover, internal states encode broader linguistic patterns than truthfulness, so it is intuitive to investigate whether integrating uncertainty can act as a source of truthfulness to amplify the internal representations. In this work, we seek to determine whether incorporating uncertainty metrics into the original token representation space can enhance their discriminative capability in identifying hallucinations.
The collection of predictive uncertainty measurements can be categorised into three levels: i) token-level uncertainty, i.e., predictive probabilities:
$P_{\text{uncertainty}}^{t}=P(y|y_{<i},\mathbf{x}),$ (7)
where $\mathbf{x}$ are prompt tokens, $y$ are generated tokens and $i\in\{1,2,\cdots,t\}$; ii) sentence-level perplexity, which is monotonically related to the mean of the negative log-likelihood of the sequence:
$P_{\text{uncertainty}}^{s}=-\frac{1}{T}\sum_{t=1}^{T}\log P(y|y_{<i},\mathbf{x}),$ (8)
and iii) semantic consistency across multiple samples, where the uncertainty value can be quantified by the number of semantic-equivalence generations over the whole generations based on the entailment results. We choose Semantic Entropy here farquhar2024detecting . Specifically,
$P_{\text{uncertainty}}^{c}=\sum_{1}^{M}\frac{I_{c}=c_{m}}{M},$ (9)
where $M$ is the total number of generations, $c_{m}$ is the identified cluster and $I_{c}$ is the assigned cluster identity of a given sample.
These uncertainty metrics can be directly leveraged for hallucination detection, and on average, the semantic consistency outperforms the other two metrics. However, it requires more computational cost while the other two metrics can obtained from just one generation as long as we can access the internal states of LLMs. To augment the internal representations with the uncertainty information, we define the final input representation for each token as follows:
$\mathbf{h}^{\prime}=(1+\lambda\cdot P_{\text{uncertainty}})\cdot\mathbf{h},$ (10)
where $P_{\text{uncertainty}}$ can be instantiated by one of the above three measurements, and $\lambda$ is used to control the impact of uncertainty metrics. The improvements gained by various uncertainty measurements and the choice of $\lambda$ are evaluated and discussed in Section 5.3.
@@ 5 Experiments
We evaluate our method on four popular benchmark datasets across a range of question-answering (QA) domains, including (1) Trivia QA joshi2017triviaqa , a relatively complicated confabulation QA datasets; (2) Stanford Question Answering Dataset (SQuAD for short) rajpurkar2016squad , which is based on Wikipedia and generated by humans through crowdsourcing; (3) Natural Questions (denoted as NQ) kwiatkowski2019natural , containing search information from real users on Google search and (4) a biomedical QA corpus BioASQ krithara2023bioasq . For each dataset, we randomly extract $2,000$ QA pairs for training and $800$ pairs for testing. For multi-sampling approaches, we prompt LLMs six times for each question for the test set generation. Following farquhar2024detecting , we have a total of $500$ questions randomly sampled from the generated test set, with approximately $400$ questions retained as the final inputs at the testing stage for evaluation. The refined set from the remaining $300$ QA pairs is used as a validation set for selecting the optimal layer for representation extraction. We employ representative open-sourced LLMs, LLaMA touvron2023llama ; dubey2024llama and Mistral family jiang2023mistral , and evaluate our approach on LLaMA-3.1-8B and Mistral-Nemo-Instruct (12B), as well as a larger version of LLaMA, namely LLaMA-3.13-Instruct-70B, which is deployed with 8-bit quantisation wolf-etal-2020-transformers . All answers are generated under the temperature of $0.5$ and with context-free zero-shot prompts. Please see more details for the setup in Appendix.
@@
We compare our method with a series of state-of-the-art (SotA) methods, covering both uncertainty-based methods and internal representation-based methods. The uncertainty-based methods include: p(True) kadavath2022language , that asks LLMs to express the correctness for given answers themselves; Perplexity ren2023outofdistribution , which is based on Eq. 8; Semantic Entropy (SE) farquhar2024detecting , semantic-equivalence measurement across multiple samples, where GPT-3.5 is used for entailment evaluation; MARS bakman-etal-2024-mars , that performs a weighted aggregation of token logits, and its enhanced version (MARS-SE), augmented with semantic entropy. For internal representation-based methods, probing classifiers with various supervision signals are included: CCS burns2022discovering , utilising contrast-consistent discovering for correctness probability measurements; SAPLMA azaria-mitchell-2023-internal , trained with correctness labels; HaloScope du2024haloscope , which proposes an automated membership estimation score and converts the score to binary labels for classification. All these methods are based on the last generated token as settled in their original paper. Notably, SAPLMA utilises a four-layer multilayer perception (MLP) with ReLU non-linearity while the other methods, including HaMI, employ a detector with two layers. We also assess the performance of HaMI using varying number of layers in Section 5.3.
Different variants of HaMI are evaluated in our experiments depending on the instantiation of Eq. 10. HaMI uses $P_{\text{uncertainty}}^{c}$ in Eq. 10 by default due to its superior performance, while HaMI is the basic variant of HaMI that does not involve Eq. 10, using solely the original features. $\lambda$ in Eq. 4 is set to 1.0 by default. Unless otherwise specified, the hidden dimension of the two-layer MLP is set to 256. $k$ in Eq. 3 is dynamically determined by the generated token length ($l$), defined as $k=\lfloor 0.1\times l\rfloor+1$.
@@
Following farquhar2024detecting ; ren2023outofdistribution , we evaluate the model’s capability for hallucination detection by calculating the area under the receiver operating characteristic curve (AUROC). The ground-truth labels indicating the correctness are given by GPT-4.1 achiam2023gpt , which is instructed to determine if the answer is correct or not based on the consistency between generated answers and gold answers and their own knowledge. The prompt template follows farquhar2024detecting (see more details in Appendix). Note that since GPT-4.1 makes mistakes mostly for the positive samples, we ask GPT-4.1 to rejudge samples labelled as positive and discard samples if the result is inconsistent with the first result. For this reason, the final sizes of the training and testing sets vary around $1,900$ and $400$, respectively.
@@ 5.1 Main Results
In Table 1, we evaluate our proposed HaMI by comparing it with eight SotA competing detection methods on four diverse QA datasets in LLMs from two different families of various sizes. As depicted in the table, our approach achieves superior performance compared with other methods in all three LLMs. In particular, HaMI consistently and substantially outperforms all methods that do not require assistance from external LLMs, including Perplexity, CCS, SAPLMA and HaloScope. When compared to approaches that leverage LLMs at the post-generation stage, HaMI achieves similar superiority across all cases except on the SQuAD dataset with LLaMA-3.1-8B, where its performance is comparable to the best score achieved by the SE method.
Moreover, with the augmentation of the uncertainty measurement, we observe that HaMI significantly outperforms all SotA methods by a large margin. It gains as large as 8.1% to 11.9% averaged AUROC improvement over MARS-SE, which also leverages entailment information among various generations, in three LLMs. Notably, p(True), MARS and Semantic Entropy all resort to external LLMs for assistance, but their capability for hallucination detection differs significantly. Following the configurations described in their respective original studies, p(True) and MARS employ relatively smaller and less capable models, whereas Semantic Entropy harnesses the more powerful GPT-3.5 model. This variation underscores that the performance of uncertainty-based methods requiring multiple sampling is highly dependent upon the capabilities of the selected assistant LLM.
Furthermore, among methods based on internal representations, supervised approaches generally outperform the semi-supervised/unsupervised ones. For example, SAPLMA and our approach exhibit superior performance to CCS and HaloScope. Notably, both HaMI and HaMI outperform SAPLMA and demonstrate consistently strong results across all four benchmarks. This superiority is due to not only the predictive uncertainty enhancement but also the effectiveness of adaptive token selection on generation responses of diverse forms (see Section 5.3 for more detailed analysis).
Comparing the performance on LLaMA models with various scales, we observe that HaMI achieves a notably larger margin of improvement compared to baselines for larger LLaMA-3.13-Instruct-70B. For instance, on the LLaMA-3.1-8B model, HaMI surpasses SE and SAPLMA by an average of 7.4% and 5.5%, respectively. On the larger model, these improvements increase to 11.5% and 8.7%, representing a relative improvement of over 50%.
LLaMA-3.1-8B Mistral-Nemo-Instruct (12B) LLaMA-3.3-Instruct-70B
Trivia QA SQuAD NQ BioASQ Trivia QA SQuAD NQ BioASQ Trivia QA SQuAD NQ BioASQ
p(True) kadavath2022language 0.666 0.614 0.650 0.673 0.862 0.749 0.772 0.749 0.563 0.596 0.530 0.573
Perplexity ren2023outofdistribution 0.732 0.649 0.659 0.709 0.720 0.662 0.646 0.675 0.671 0.626 0.619 0.583
SE farquhar2024detecting 0.828 0.787 0.773 0.757 0.795 0.752 0.813 0.800 0.819 0.643 0.769 0.772
MARS bakman-etal-2024-mars 0.766 0.663 0.661 0.706 0.765 0.693 0.700 0.710 0.704 0.648 0.635 0.610
MARS-SE bakman-etal-2024-mars 0.824 0.780 0.777 0.744 0.797 0.743 0.786 0.798 0.794 0.676 0.772 0.769
CCS burns2022discovering 0.675 0.596 0.628 0.662 0.551 0.579 0.615 0.553 0.597 0.575 0.592 0.562
SAPLMA azaria-mitchell-2023-internal 0.835 0.769 0.781 0.821 0.860 0.797 0.836 0.859 0.842 0.672 0.817 0.748
HaloScope du2024haloscope 0.610 0.669 0.637 0.635 0.653 0.621 0.568 0.676 0.656 0.655 0.622 0.556
HaMI (Ours) 0.854 0.783 0.788 0.823 0.883 0.813 0.843 0.873 0.858 0.765 0.820 0.813
HaMI (Ours) 0.897 0.826 0.820 0.836 0.903 0.837 0.867 0.888 0.891 0.774 0.846 0.825
Table 1: AUROC results of HaMI and its competing methods on four datasets with three LLMs. HaMI denotes a basic HaMI using the original representations as inputs. The best results are in red and second-best are in blue.
@@
Figure 3: (a) AUROC results of cross-dataset generalisation on four datasets using LLaMA-3.1-8B. (b) AUROC w.r.t. dimensionality of the feature layer (left) and the number of network layers (right) based on LLaMA-3.1-8B.
@@ 5.2 Cross-dataset Generalisation Ability
The ability to generalise across datasets is essential to facilitating real-world applications of LLMs in diverse domains. We conduct experiments on the aforementioned four datasets to assess whether the proposed HaMI can effectively generalise among various datasets, with three strong competing methods as baselines. For each dataset, we report the average AUROC scores of detectors trained on one of the other three datasets. The results are illustrated in Figure 3(a). It is clear that our method HaMI achieves consistently the best generalisation performance across all four datasets, outperforming the best competing method SAPLMA by 7% - 9%. Compared to the within-dataset performance in Table 1, the maximum performance decline in HaMI is no more than $4.5\%$, observed on the BioASQ dataset, which is significantly lower than that of the competing methods. These findings affirm the effectiveness of HaMI as a reliable hallucination detector in unseen datasets.
@@ 5.3 Ablation Study
Figure 4: Adaptive token selection results showing tokens with the top-2 highest hallucination scores.
LLaMA-3.1-8B Mistral
Trivia QA SQuAD Trivia QA SQuAD
First 0.849 0.774 0.844 0.799
Before Last 0.878 0.778 0.849 0.807
Last 0.890 0.804 0.873 0.825
Ours 0.897 0.826 0.903 0.837
LLaMA-3.1-8B Mistral
Trivia QA SQuAD Trivia QA SQuAD
Original 0.854 0.783 0.883 0.813
$P_{\text{uncertainty}}^{t}$ 0.856 0.782 0.884 0.825
$P_{\text{uncertainty}}^{s}$ 0.871 0.787 0.897 0.831
$P_{\text{uncertainty}}^{c}$ 0.897 0.826 0.903 0.837
Table 2: Ablation study results on the ATS module (left) and uncertainty-enabled representation module (right).
Analysis of MIL-based Adaptive Token Selection.
Unlike existing methods that predetermined critical tokens for capturing truthfulness information, HaMI proposes the MIL-based adaptive token selection (ATS) module. Table 2 Left compares the performance of our ATS module with commonly used methods, including the First generated token, the Last generated token, and the Before Last generated token. The results show that our ATS method outperforms the alternatives across both Trivia QA and SQuAD datasets on both LLMs, yielding an average improvement of 6% over the First token, 4% over the Before Last token and 2% over the Last token.
Furthermore, our ATS module exhibits substantially better robustness to the length of generations than the competing methods. In particular, the average number of generated tokens for the Trivia QA dataset with the LLaMA model is $10$, which is smaller than the other three cases (Trivia QA with Mistral model - $13.5$, SQuAD with LLaMA model - $17.5$, SQuAD with Mistral model - $15$). Despite this variation, the ATS module performs robustly on all four cases and even performs more effectively as the length increases. In contrast, the performance of the predefined token location methods performs unstably and tends to degrade as the length of the generation grows. Moreover, the Last token achieves better performance across all scenarios, and the First token is just the opposite. It suggests that the last token can capture full semantics under specific length conditions, while the first token may miss content in the subsequent predictions.
@@
Figure 4 presents an illustrative example of the scoring results over tokens in a positive bag and a negative bag, where we can observe that there is significant distinguishability between the positive and negative tokens since the maximum scores of the instances in the positive bag is substantially greater than that in the negative bag. Tokens denoted by blue characters have scores that are very close to the largest ones and we find that these tokens can be adaptively identified by our method as concentrated answers in comparison to ground-truth answers. Additionally, it is noted that while selected tokens are associated with the exact answer, they may appear at any location in their vicinity, which can be captured by our smoothness loss.
Analysis of Different Predictive Uncertainty-enabled Features.
@@
Here we systematically examine three uncertainty measuring methods presented in Section 4.2 for the internal representation enhancement as detailed in Eq. 7, 8 and 9 respectively. We use a simple combination strategy as defined in Eq. 10. In the right panel of Table 2, we can observe that both $P_{\text{uncertainty}}^{s}$ and $P_{\text{uncertainty}}^{c}$ contribute to improved detection capabilities over the baseline (i.e.,, the original feature representations) while the token-level uncertainty $P_{\text{uncertainty}}^{t}$ yields limited effectiveness. Specifically, the enhancements attributed to semantic consistency $P_{\text{uncertainty}}^{c}$ are the most significant, exhibiting improvements up to $6.7\%$. These observations suggest that the effectiveness of uncertainty-based enhancement is tied to the inherent ability of the uncertainty metric to distinguish hallucinations from correct responses. Notably, although the improvements observed with $P_{\text{uncertainty}}^{s}$ are less pronounced, it surpasses the performance of SotA multi-sampling approaches (such as SE in Table 1), without incurring the costs associated with multiple generations and external LLM employments. This highlights the potential of HaMI for deployments in various practical environments that involve external tools or not.
Figure 5: (a) Performance of HaMI w.r.t. internal representations extracted from different layers. (b) The impact of augmentation strength $\lambda$ in Eq. 10 on HaMI. All results are based on LLaMA-3.1-8B.
Performance of HaMI w.r.t. Representations from Different LLM Layers.
@@
We evaluate detection performance using representations extracted from various layers of LLM with LLaMA-3.1-8B across all benchmarks. We apply both original internal representations and semantic-equivalence-enhanced representations for investigation. As depicted by the dashed line in Figure 5(a), the AUROC values for original representations exhibit a clear increase, peaking between layers 12 and 16, before declining to a relatively stable level. This observation suggests that the truthfulness content evolves across the initial to middle layers. The performance of the uncertainty-enhanced representations (presented in the solid line) maintains a relatively consistent trend, with the highest AUROC scores concentrated in the middle layers as well. The comparison of these trends indicates that incorporating predictive uncertainty enhances the distinctiveness of the representations, particularly in the earlier layers where less semantic information is typically available.
Performance of HaMI w.r.t. Various Model Scales.
We investigate the sensitivity of HaMI on the model scale by varying the number of layers and hidden dimensions in the two-layer MLP network. As illustrated in Figure 3(b), HaMI exhibits low sensitivity to both hyperparameters across all four benchmarks. Specifically, for a two-layer MLP (employed in HaMI and most baselines), the performance remains steady as the size of the hidden dimension changes from 64 to 512. Similarly, increasing the number of layers from 1 to 4 yields only marginal changes in AUROC, suggesting that a shallow architecture is sufficient for effective hallucination detection. These findings highlight the robustness of our method w.r.t. variations in network architecture.
Performance of HaMI w.r.t. Uncertainty Augmentation Strength ($\lambda$).
Figure 5(b) presents results based on various choice of $\lambda$ used in Eq. 10. It is clear that the model’s performance is steadily improved as $\lambda$ changed from 0.0 towards 1.0. Beyond this point, performance becomes stable.